# From Selection to Generation: A Survey of LLM-based Active Learning

### v1

@@ 5.2 Cost-Aware Termination
In LLM-based AL, estimating the annotation cost can be challenging and complex. While traditional AL relies on a discrete budget $k$ to represent the number of human annotations, LLM-based AL may combine both human and LLM annotations. In such cases, the cost of an annotation depends not only on the source (human vs. LLM) but also on variable factors like input and output token counts. Even in the case that LLM-based annotations are used without any human annotations, the cost cannot be easily approximated as the discrete budget that often represents the amount of examples that can be labeled. It is straightforward to see that since the cost depends on the input and more so the output tokens, then the budget may be better represented as a real-valued amount that pertains directly to a monetary cost. Recent exploration in hybrid termination criteria Akins et al. (2024); Pullar-Strecker et al. (2024) including cost-aware termination criterion that integrates token-level cost analysis and performance plateau detection. Other works have suggested a combined cost-performance metric that balances the fixed costs of human annotations with the variable costs of LLM-based annotations Xia et al. (2024a); Zhang et al. (2024a).
@@ 6 AL Paradigms with LLMs
Active Supervised Finetuning
To reduce labeling costs, active learning has been integrated into supervised finetuning via uncertainty-based querying, self-training on low-uncertainty data, and strategic sample selection (Yu et al., 2022; Xia et al., 2024b; Bayer and Reuter, 2024).

### v2

@@ 5.2 Cost-Aware Termination
In LLM-based AL, estimating the annotation cost can be challenging and complex. While traditional AL relies on a discrete budget $k$ to represent the number of human annotations, LLM-based AL may combine both human and LLM annotations. In such cases, the cost of an annotation depends not only on the source (human vs. LLM) but also on variable factors like input and output token counts. Even in the case that LLM-based annotations are used without any human annotations, the cost cannot be easily approximated as the discrete budget that often represents the amount of examples that can be labeled. It is straightforward to see that since the cost depends on the input and more so the output tokens, then the budget may be better represented as a real-valued amount that pertains directly to a monetary cost. Recent exploration in hybrid stopping criteria Akins et al. (2024); Pullar-Strecker et al. (2024) including cost-aware termination criterion that integrates token-level cost analysis and performance plateau detection. Other works have suggested a combined cost-performance metric that balances the fixed costs of human annotations with the variable costs of LLM-based annotations Xia et al. (2024a); Zhang et al. (2024a).
@@ 6 AL Paradigms with LLMs
Active Supervised Fine-Tuning
To reduce labeling costs, active learning has been integrated into supervised fine-tuning via uncertainty-based querying, self-training on low-uncertainty data, and strategic sample selection (Yu et al., 2022; Xia et al., 2024b; Bayer and Reuter, 2024).
@@ 8 Open Problems & Challenges
Complex Feedback and Multi-Aspect Labels
Traditional AL usually relies on single-label supervision, but reward models in the era of LLMs are able to return structured feedback that scores multiple facets of a response Bai et al. (2022), such as factual accuracy, completeness, and fluency. Developing LLM-based AL strategies that can leverage these rich signals without overfitting to any one dimension may be a promising solution.